OpenStudy (anonymous):

Word Problem help please 1. The magician said "The average of seven numbers is 49. If 1 is added to the fisr number, 2 is added to the second number, 3 is added to the third number and so on up to seventh number". what is the new average ?

OpenStudy (anonymous):

can u help

OpenStudy (ksaimouli):

11 is the average

OpenStudy (anonymous):

step by step because i have no idea how u got that

OpenStudy (anonymous):

\[49=\frac{x_1+x_2+...+x_7}{7}\] \[\frac{x_1+1+x_2+2+..+x_7+7}{7}=\frac{x_1+x_2+...+x_7+28}{7}\] \[=49+\frac{28}{7}=49+4=53\]

OpenStudy (anonymous):

oh kk i have another word problem could u people help me

OpenStudy (anonymous):

snaap ...

jimthompson5910 (jim_thompson5910):

The average is equal to the sum of all the numbers divided by the number of numbers. So if the numbers add up to some number S, then the average A is A = S/7 But we know that the average is 49, so A = 49, which means that A = S/7 49 = S/7 49*7 = S 343 = S So S = 343 Since we're adding 1 to the first, 2 to the second, etc all the way up to 7 to the seventh, we're basically adding 1+2+3+4+5+6+7 = 28 to S to get S+28, which means that the new sum is now S+28 = 343 + 28 = 371 Now divide this new sum by 7 to get 53 So the new average is 53

OpenStudy (anonymous):

wait im getting confused ?

OpenStudy (anonymous):

(49*7+28)/7=53

OpenStudy (anonymous):

\[x_1+1+x_2+2+x_3+3+...+x_7+7 \] \[(x_1+x_2+..+x_7)+(1+2+3+4+5+6+7)\]

OpenStudy (anonymous):

whats the real answer

OpenStudy (anonymous):

11 or 53

OpenStudy (anonymous):

53

OpenStudy (anonymous):

satellite and jim's 53

OpenStudy (anonymous):

kk can u help with another one

OpenStudy (anonymous):

kk can u help with another one

OpenStudy (anonymous):

anybody

OpenStudy (anonymous):

You can post a new question so all can see it.

OpenStudy (anonymous):

kk in a given school, there 240 boys and 260 girls ...what is the ratio of numbers of girls to the number of boys ...what is the ratio of the number of boys to the total number of pupils in the school?

OpenStudy (anonymous):

ratio of girls to boys = 260 / 240

OpenStudy (anonymous):

ratio of boys to total pupils = 240 / (240+260) = 240 / 500

OpenStudy (anonymous):

?? for first ? what isnt that just telling me girls to boys 260/240??

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Ratio

OpenStudy (anonymous):

ok so ready for next ?

OpenStudy (anonymous):

why don't you ask a new question ?

OpenStudy (anonymous):

1. if tim had lunch at $50.50 and he gave 20% tip. how much did he spend? 2. Find the greatest common factor of 24, 40 and 60 ? 3. in 2008 , the world population is about 6,760,000,000. Write the 2008 wolrd population in scientific notation 4. give the rules of divisibility ?

OpenStudy (anonymous):

some of the rules